# AVadCLIP: Audio-Visual Collaboration for Robust Video Anomaly Detection

### v1

@@ Abstract
With the increasing adoption of video anomaly detection in intelligent surveillance domains, conventional visual-based detection approaches often struggle with information insufficiency and high false-positive rates in complex environments. To address these limitations, we present a novel weakly supervised framework that leverages audio-visual collaboration for robust video anomaly detection. Capitalizing on the exceptional cross-modal representation learning capabilities of Contrastive Language-Image Pretraining (CLIP) across visual, audio, and textual domains, our framework introduces two major innovations: an efficient audio-visual fusion that enables adaptive cross-modal integration through lightweight parametric adaptation while maintaining the frozen CLIP backbone, and a novel audio-visual prompt that dynamically enhances text embeddings with key multimodal information based on the semantic correlation between audio-visual features and textual labels, significantly improving CLIP’s generalization for the video anomaly detection task. Moreover, to enhance robustness against modality deficiency during inference, we further develop an uncertainty-driven feature distillation module that synthesizes audio-visual representations from visual-only inputs. This module employs uncertainty modeling based on the diversity of audio-visual features to dynamically emphasize challenging features during the distillation process. Our framework demonstrates superior performance across multiple benchmarks, with audio integration significantly boosting anomaly detection accuracy in various scenarios. Notably, with unimodal data enhanced by uncertainty-driven distillation, our approach consistently outperforms current unimodal VAD methods.
1 Introduction
Video anomaly detection (VAD), as a pivotal technology in intelligent surveillance systems, focuses on identifying anomalous events within videos and has attracted substantial research interest in recent years [30, 13, 47]. Due to the rarity of anomalies and the high cost of manual annotation, fully supervised frameworks are impractical for large-scale deployment. As a solution, weakly supervised video anomaly detection (WSVAD) [38, 58, 18, 56] has gained traction, requiring only video-level labels instead of detailed frame annotations. These methods seek to uncover latent anomalies under limited supervision, addressing challenges such as anomaly insufficiency and high false-positive rates in complex environments.
@@ 1 Introduction
However, these methods—whether CLIP-based or conventional—predominantly rely on unimodal visual information, which often leads to significant detection limitations in complex real-world scenarios. Visual occlusion, extreme lighting variations, and environmental noise can render visual features unreliable or ambiguous [46, 39, 40]. In these challenging conditions, multimodal information, particularly audio, offers indispensable contextual cues that can complement and enhance visual-based detection. For instance, audio remains robust when visual data is compromised, allowing detection of off-camera events. In acoustically rich environments, certain anomalies like explosion, scream, or gunshot exhibit distinct acoustic signatures, making them more discriminative in the audio domain. Similarly, in low-light conditions where visual features degrade, audio serves as a critical supplementary modality. These observations underscore the importance of integrating audio and video modalities, as their complementary nature can significantly enhance the accuracy and robustness of anomaly detection systems in diverse and challenging environments. We illustrate the impact of audio-visual integration for WSVAD in Figure 1.
Existing attempts [45, 46, 62] to incorporate audio into video anomaly detection typically adopt traditional feature concatenation methods, such as fusing visual features extracted by I3D [2] or C3D [42] with audio features extracted by VGGish [12]. These approaches fail to fully exploit the potential of modern multimodal learning techniques, resulting in suboptimal cross-modal integration. Moreover, they overlook the inherent semantic correlations between visual and auditory modalities, which are essential for enhancing anomaly detection performance.
@@
$\bullet$ We propose a weakly supervised video anomaly detection framework that harnesses audio-visual collaborative learning, leveraging CLIP’s multimodal alignment capabilities. By incorporating a lightweight adaptive audio-visual fusion mechanism and integrating audio-visual information through prompt-based learning, our approach effectively achieves CLIP-driven robust anomaly detection in multimodal settings.
@@
$\bullet$ We design an uncertainty-driven feature distillation module, which transforms deterministic estimation into probabilistic uncertainty estimation. This enables the model to capture feature distribution variance, ensuring robust anomaly detection performance even with unimodal data.
@@
$\bullet$ Extensive experiments on two WSVAD datasets demonstrate that our method achieves superior performance in audio-visual scenarios, while maintaining robust anomaly detection results even in audio-absent conditions.
2 Related Work
2.1 Video Anomaly Detection
@@
Video Anomaly Detection has been extensively studied in recent years, with existing approaches broadly categorized into semi-supervised and weakly supervised methods. Among them, semi-supervised methods primarily rely on normal video clips for training and identify anomalies by detecting deviations from learned normal patterns during inference. These methods commonly adopt self-supervised learning techniques [19, 37, 20], such as reconstruction [10, 29] or prediction [27, 1]. Reconstruction-based methods assume that the model can effectively reconstruct normal videos, whereas abnormal videos—due to distributional discrepancies—result in significant reconstruction errors. Autoencoders [16, 14] are widely employed to capture normal pattern features, with reconstruction error serving as an anomaly indicator. Prediction-based methods [55] utilize models to forecast future frames, detecting anomalies based on prediction errors. However, a key limitation of semi-supervised methods is their tendency to overfit normal patterns, leading to poor generalization to unseen anomalies.
@@
Weakly supervised methods, in contrast, typically adopt the MIL framework, requiring only video-level anomaly labels and significantly reducing annotation costs. A classic work is DeepMIL [38], which employs a ranking loss to distinguish normal from anomalous instances. To enhance detection, two-stage pseudo-labeling strategy has been introduced, where high-confidence anomalous regions identified during MIL training serve as pseudo-labels for a secondary refinement phase [25, 11, 9]. With the rise of Vision-Language Models (VLMs) [5], CLIP has shown remarkable cross-modal capabilities and is increasingly applied to WSVAD. VadCLIP [50], the first CLIP-based WSVAD method, integrates textual priors via text and visual prompts, enhancing anomaly detection. Building on this, TPWNG [56] refines feature learning through a two-stage approach. Recent research trends focus on large model-driven strategies, e.g., training-free frameworks [59, 53], spatiotemporal anomaly detection [49], and open-scene anomaly detection [48].
2.2 Audio-Visual Learning
@@
3 Methodology
Figure 2: The pipeline of our proposed AVadCLIP.
Figure 3: The pipeline of our proposed adaptive fusion module and binary classifier.
3.1 Problem Statement
@@
3.2 Video and Audio Encoders
@@
3.3 Audio-Visual Adaptive Fusion
@@
3.4 Dual Branch Framework with Prompts
We leverage a two-branch framework [50] for the WSVAD task, consisting of a classification branch and an alignment branch, which effectively leverage audio-visual information to improve detection accuracy.
Classification branch consists of a lightweight binary classifier (as shown in Figure 3), which takes $X_{av}$ as input and directly predicts the frame-level anomaly confidence $A$.
Alignment branch leverages the cross-modal semantic alignment mechanism, which computes the similarity between frame-level features and class label features. To obtain class label representations, we leverage the CLIP text encoder combined with the learnable textual prompt [65] and audio-visual prompt to extract class embeddings, ensuring unified semantic alignment between visual and textual modalities. Given a set of predefined class labels (e.g., “normal” “fighting”), we first introduce a learnable textual prompt, then we concatenate the textual prompt with class labels and feed them into CLIP text encoder to obtain the class representation $X_{c}$. Compared to the manually defined prompt, the learnable prompt allows the model to dynamically adjust textual representations during training, making them more suitable for the specific requirements of WSVAD. Furthermore, we incorporate an audio-visual prompt into the class label features to enrich the class representations with additional multimodal information.
@@ 3.4 Dual Branch Framework with Prompts
This two-branch framework provides anomaly confidence through the classification branch and refines category identification with class information via the alignment branch, improving robustness and enabling fine-grained anomaly detection.
3.5 Optimization of Audio-Visual Model
@@
3.6 Uncertainty-Driven Distillation
In the WSVAD task, audio serves as a complementary modality to video, enhancing detection accuracy. However, audio may be unavailable in practical scenarios, leading to performance degradation. To address this, we apply knowledge distillation (KD) by using a pre-trained multi-modal (video+audio) teacher model to guide a unimodal (video-only) student model, ensuring robust anomaly detection even without audio. Traditional knowledge distillation methods typically assume a deterministic transfer of knowledge, employing mean square error (MSE) loss to align the student model with the teacher’s feature representations. However, this approach fails to account for the inherent uncertainty in audio-visual feature fusion. In real-world scenarios, factors such as noisy audio or occluded visual content can introduce distortions in the fused features, leading to inaccurate feature representations and diminished generalization capability.
To overcome this, we propose a probabilistic uncertainty distillation strategy [4, 54], which models data uncertainty during distillation, improving the student model’s robustness across diverse scenarios. Specifically, assume $X_{av,i}=X_{vs,i}+\epsilon\sigma_{i}$, where $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, $X_{vs}$ represents enhanced visual features derived from the student model, and $\sigma_{i}$ refers to the inherent uncertainty between the $i^{th}$ pair of features. Then we model the observation as a Gaussian likelihood function to more accurately quantify data uncertainty in the feature distillation. The relationship between the audio-visual fusion feature $X_{av,i}$ and the unimodal feature $X_{vs,i}$ is formulated as:
@@ 3.6 Uncertainty-Driven Distillation
In practice, we design a network branch (a simple three-layer convolutional neural network) to predict the variance ${\sigma}_{i}^{2}$ and reformulate the likelihood maximization problem as the minimization of a loss function. Specifically, we employ an uncertainty-weighted MSE loss:
@@
4 Experiments
4.1 Datasets and Evaluation
@@
XD-Violence. As the largest publicly available audio-visual WSVAD dataset, XD-Violence [45] significantly surpasses existing datasets in scale and diversity. It comprises 3,954 training videos and 800 test videos, with the test set containing 500 violent and 300 non-violent instances. The dataset covers six distinct categories of violent events, including abuse, car accidents, explosion, fighting, riot, and shooting, which occur at various temporal locations within videos.
@@
CCTV-Fightssub. Derived from CCTV-Fights [35], CCTV-Fightssub [32] is a carefully curated subset designed to address audio-visual anomaly detection, The subset retains 644 high-quality videos depicting real-world fight scenarios, each with meaningful audio content, making it a valuable resource for evaluating audio-visual anomaly detection methods in real-world surveillance contexts.
@@
Evaluation. For performance evaluation, we adopt distinct metrics tailored to different granularities of WSVAD tasks. For coarse-grained WSVAD, we employ frame-level Average Precision (AP), which provides a comprehensive measure of detection accuracy across varying confidence thresholds. For fine-grained anomaly detection, we utilize mean Average Precision (mAP) [46] computed across multiple intersection over union (IoU) thresholds and the average mAP (AVG) across different thresholds.
4.2 Implementation Details
@@
We conduct experiments on an NVIDIA RTX 4090 GPU, where the visual enhancement module is a single-layer 1D convolutional network, which includes a convolutional layer with a kernel size of 3, padding size of 1, ReLU activation function, and a skip connection. Such an operation effectively facilitates the aggregation of local contextual information. For input processing, we employ a frame selection strategy tailored to different datasets, sampling one frame per 16 frames for XD-Violence and one frame per 4 frames for CCTV-Fightssub, using a uniform sampling strategy with a maximum frame count of 256; During optimization, we set the batch size, learning rate, and total epoch to 96, $1e^{-5}$, and 10, respectively.
4.3 Comparison with State-of-the-Art Methods
Performance comparison on XD-Violence
Our experiments evaluate both coarse-grained and fine-grained anomaly detection performance on XD-Violence, comparing our AVadCLIP against state-of-the-art approaches, as shown in Tables 1 and 2.
For coarse-grained anomaly detection, using only RGB inputs, AVadCLIP (* denotes RGB-only input) achieves an AP score of 85.53%, surpassing all existing vision-only methods. Notably, it outperforms VadCLIP—the previous best-performing RGB-only approach—by 1.0%, demonstrating superior visual anomaly detection. When incorporating audio, AVadCLIP further improves performance, significantly outperforming all multimodal baselines, achieving a remarkable 4.9% gain over the latest AVCL [32] method.
For fine-grained anomaly detection, AVadCLIP consistently outperforms all competitors across different IoU thresholds, as detailed in Table 2. With RGB-only input, AVadCLIP surpasses VadCLIP at all IoU thresholds, achieving an AVG improvement of 2.7%. Similarly, the full-modality model AVadCLIP leads across all metrics, boosting the AVG by 3.9%. These results highlight the effectiveness of multimodal learning in precisely localizing anomaly boundaries and improving category predictions.
Overall, AVadCLIP achieves state-of-the-art performance in both unimodal and multimodal settings across coarse- and fine-grained anomaly detection tasks. The comprehensive results validate its effectiveness in leveraging audio-visual collaboration and demonstrate the feasibility of uncertainty-driven distillation strategy.
@@ 4.3 Comparison with State-of-the-Art Methods
Table 1: Coarse-grained comparisons on XD-Violence.
Method mAP@IoU(%)
0.1 0.2 0.3 0.4 0.5 AVG
Random 1.82 0.92 0.48 0.23 0.09 0.71
DeepMIL [38] 22.72 15.57 9.98 6.20 3.78 11.65
AVVD [46] 30.51 25.75 20.18 14.83 9.79 20.21
VadCLIP [50] 37.03 30.84 23.38 17.90 14.31 24.70
AVadCLIP 39.63 32.77 26.84 21.58 16.39 27.44
AVadCLIP 41.89 34.61 27.08 22.16 17.30 28.61
Table 2: Fine-grained comparisons on XD-Violence.
@@
Table 3: Coarse-grained comparisons on CCTV-Fightssub.
Performance comparison on CCTV-Fightssub
The coarse-grained anomaly detection results on CCTV-Fightssub are presented in Table 3. For RGB-only methods, AVadCLIP achieves 73.36% AP, surpassing the state-of-the-art VadCLIP and demonstrating the effectiveness of our approach in unimodal scenarios. For audio-visual scenarios, AVadCLIP further improves performance, outperforming all existing methods. These results indicate that incorporating audio information can further enhance anomaly detection performance, validating the effectiveness of cross-modal complementary information mining.
4.4 Ablation Studies

### v2

@@ Abstract
With the increasing adoption of video anomaly detection in intelligent surveillance domains, conventional visual-only detection approaches often struggle with information insufficiency and high false-positive rates in complex environments. To address these limitations, we present a novel weakly supervised framework that leverages audio-visual collaboration for robust video anomaly detection. Capitalizing on the exceptional cross-modal representation learning capabilities of Contrastive Language-Image Pretraining (CLIP) across visual, audio, and textual domains, our framework introduces two major innovations: an efficient audio-visual fusion that enables adaptive cross-modal integration through lightweight parametric adaptation while maintaining the frozen CLIP backbone, and a novel audio-visual prompt that dynamically enhances text embeddings with key multimodal information based on the semantic correlation between audio-visual features and textual labels, significantly improving CLIP’s generalization for the video anomaly detection task. Moreover, to enhance robustness against modality deficiency during inference, we further develop an uncertainty-driven feature distillation module that synthesizes audio-visual representations from visual-only inputs. This module employs uncertainty modeling based on the diversity of audio-visual features to dynamically emphasize challenging features during the distillation process. Our framework demonstrates superior performance across multiple benchmarks, with audio integration significantly boosting anomaly detection accuracy in various scenarios. Notably, with unimodal data enhanced by uncertainty-driven distillation, our approach consistently outperforms current unimodal VAD methods.
Index Terms:
video anomaly detection, audio-visual collaboration, weakly supervised learning.
I Introduction
Video anomaly detection (VAD), as a pivotal technology in intelligent surveillance systems, focuses on identifying anomalous events within videos and has attracted substantial research interest in recent years [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. Due to the rarity of anomalies and the high cost of manual annotation, fully supervised frameworks are impractical for large-scale deployment. As a solution, weakly supervised video anomaly detection (WSVAD) methods [12, 13, 14, 15] have gained traction, aiming to discover latent anomalies under coarse supervision.
@@ I Introduction
However, these methods, whether CLIP-based or conventional, predominantly rely on unimodal visual information, which often leads to significant detection limitations in complex real-world scenarios. Visual occlusion, extreme lighting variations, and environmental noise can render visual features unreliable or ambiguous [22, 23, 24]. In these challenging conditions, multimodal information, particularly audio, offers indispensable contextual cues that can complement and enhance visual-based detection. For instance, audio remains robust when visual data is compromised, allowing detection of off-camera events. In acoustically rich environments, certain anomalies like explosion, scream, or gunshot exhibit distinct acoustic signatures, making them more discriminative in the audio domain. Similarly, in low-light conditions where visual features degrade, audio serves as a critical supplementary modality. These observations underscore the importance of integrating audio and video modalities, as their complementary nature can significantly enhance the accuracy and robustness of anomaly detection systems in diverse and challenging environments. We illustrate the impact of audio-visual integration for WSVAD in Figure 1.
Existing attempts [25, 22, 26] to incorporate audio into video anomaly detection typically adopt traditional feature concatenation methods, such as fusing visual features extracted by I3D [27] or C3D [28] with audio features extracted by VGGish [29]. These approaches fail to fully exploit the potential of multimodal learning, resulting in suboptimal cross-modal integration. Moreover, they overlook the inherent semantic alignment between visual and auditory modalities, which are essential for enhancing anomaly detection performance.
@@
  • We propose a WSVAD framework that harnesses audio-visual collaborative learning, leveraging CLIP’s multimodal alignment capabilities. By incorporating a lightweight adaptive audio-visual fusion mechanism and integrating audio-visual information through prompt-based learning, our approach effectively achieves CLIP-driven robust anomaly detection in multimodal settings.
@@
  • We design an uncertainty-driven feature distillation module, which transforms deterministic estimation into probabilistic uncertainty estimation. This enables the model to capture feature distribution variance, ensuring robust anomaly detection performance even with unimodal data.
@@
  • Extensive experiments on two WSVAD datasets demonstrate that our method achieves superior performance in audio-visual scenarios, while maintaining robust anomaly detection results even in audio-absent conditions.
II Related Work
II-A Video Anomaly Detection
@@
Video anomaly detection has been extensively studied in recent years, with existing approaches broadly categorized into semi-supervised and weakly supervised methods. Among them, semi-supervised methods primarily rely on normal video clips for training and identify anomalies by detecting deviations from learned normal patterns during inference. These methods commonly adopt self-supervised learning techniques [30, 31, 32], such as reconstruction [33, 34] or prediction [35, 36]. Reconstruction-based methods assume that the model can effectively reconstruct normal videos, whereas abnormal videos, due to distributional discrepancies, result in significant reconstruction errors. Autoencoders [37, 38] are widely employed to capture normal pattern features, with reconstruction error serving as an anomaly indicator. Prediction-based methods [39] utilize models to forecast future frames, detecting anomalies based on prediction errors. However, a key limitation of semi-supervised methods is their tendency to overfit normal patterns, leading to poor generalization to unseen anomalies.
@@
Weakly supervised methods, in contrast, typically adopt the MIL framework, requiring only video-level anomaly labels and significantly reducing annotation costs. The classic work, DeepMIL [12], which employs a ranking loss to distinguish normal from anomalous instances. Furthermore, two-stage self-training strategy has been proposed to further enhance detection, where high-confidence anomalous regions identified during MIL training serve as pseudo-labels for a secondary refinement phase [40, 41, 42]. With the rise of Vision-Language Models (VLMs) [43], CLIP has shown remarkable cross-modal capabilities and is increasingly applied to WSVAD. VadCLIP [21], the first CLIP-based WSVAD method, integrates textual priors via text and visual prompts, enhancing anomaly detection. Building on this, TPWNG [15] refines feature learning through a two-stage approach. Recent research trends focus on large model driven strategies, e.g., training-free frameworks [44, 45], spatiotemporal anomaly detection [46], and open-scene anomaly detection [47]. Recent advances in multi-modal fusion [48] introduce powerful frameworks combining diverse modalities such as visual and audio features. For instance, AVCL [49] and DSRL [50] have shown significant promise in improving anomaly detection by leveraging both visual and audio cues.
II-B Audio-Visual Learning
@@
II-C Large Models in Video Understanding
In recent years, large models have exhibited exceptional capabilities in perception and reasoning for video understanding tasks, significantly accelerating the shift from purely visual models to multimodal video understanding frameworks. Representative visual models, such as VideoMAE [64], employ masked self-supervised learning to effectively model spatiotemporal dynamics in videos, facilitating their widespread application in video classification, action recognition, and anomaly detection. With the success of VLMs like CLIP [18] and ALIGN [65], integrating language priors into video understanding has emerged as a prominent research trend. These models perform cross-modal semantic alignment through joint image-text encoding and have been widely adopted in tasks such as zero-shot action recognition, video retrieval, and open-vocabulary scene understanding. Further advances, including X-CLIP [66] and VideoCLIP [20], introduce temporal modeling into VLM architectures, significantly improving semantic comprehension of long-form video content. Meanwhile, VLM-based video reasoning tasks are gaining increasing attention. Models such as VL-T5 [67] and VideoChat [68] leverage language-guided mechanisms to enable video question answering, event interpretation, and causal reasoning, thereby substantially broadening the scope of video understanding.
III Methodology
Figure 2: The pipeline of our proposed AVadCLIP. Our method supports both multimodal inputs and visual-only inputs via distillation, enabling robust video anomaly detection through the proposed UKD strategy. Throughout the entire framework, the pre-trained CLIP backbone remains fully frozen, with only a few modules being trainable. This design allows for efficient and lightweight adaptation of CLIP’s knowledge to the specific task of audio-visual anomaly detection.
Figure 3: The pipeline of our proposed adaptive fusion module, binary classifier, visual enhancement network, and uncertainly modeling network.
III-A Problem Statement
@@
III-B Video and Audio Encoders
@@
III-C Audio-Visual Adaptive Fusion
@@
III-D Dual Branch Framework with Prompts
We leverage a dual-branch framework [21] for the WSVAD task, consisting of a classification branch and an alignment branch, which effectively leverage audio-visual information to improve detection accuracy. Classification branch consists of a lightweight binary classifier (as shown in Figure 3), which takes $X_{av}$ as input and directly predicts the frame-level anomaly confidence $A$. Alignment branch leverages the cross-modal semantic alignment mechanism, which computes the similarity between frame-level features and class label features. To obtain class label representations, we leverage the CLIP text encoder combined with the learnable textual prompt [73] and audio-visual prompt to extract class embeddings, ensuring unified semantic alignment between visual and textual modalities. Given a set of predefined class labels (e.g., “normal”, “fighting”), we first introduce a learnable textual prompt, then we concatenate the textual prompt with class labels and feed them into CLIP text encoder to obtain the class representation $X_{c}$. Compared to the manually defined prompt, the learnable prompt allows the model to dynamically adjust textual representations during training, making them more suitable for the specific requirements of WSVAD. Furthermore, we incorporate an audio-visual prompt into the class label features to enrich the class representations with additional multimodal information.
@@ III-D Dual Branch Framework with Prompts
This dual-branch framework provides anomaly confidence through the classification branch and refines category identification with class information via the alignment branch, improving robustness and enabling fine-grained anomaly detection.
III-E Optimization of Audio-Visual Model
@@
III-F Uncertainty-Driven Distillation
In the WSVAD task, audio serves as a complementary modality to video, enhancing detection accuracy. However, audio may be unavailable in practical scenarios, leading to performance degradation. To address this, we apply knowledge distillation by using a pre-trained multi-modal (video+audio) teacher model to guide a unimodal (video-only) student model, ensuring robust anomaly detection even without audio. Traditional knowledge distillation methods typically assume a deterministic transfer of knowledge, employing mean square error (MSE) loss to align the student model with the teacher’s feature representations. However, this approach fails to account for the inherent uncertainty in audio-visual feature fusion. In real-world scenarios, factors such as noisy audio or occluded visual content can introduce distortions in the fused features, leading to inaccurate feature representations and diminished generalization capability.
To overcome this, we propose a probabilistic uncertainty distillation strategy [75, 76], which models data uncertainty during distillation, improving the student model’s robustness across diverse scenarios. Specifically, assume $X_{av,i}=X_{vs,i}+\epsilon\sigma_{i}$, where $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, $X_{vs}$ represents enhanced visual features generated from the student model, and it is derived from $X_{v}$ after passing through a visual enhancement network, which is illustrated in Figure 3. Besides, $\sigma_{i}$ refers to the inherent uncertainty between the $i^{th}$ pair of features. Then we model the observation as a Gaussian likelihood function to more accurately quantify data uncertainty in the feature distillation. The relationship between the audio-visual fusion feature $X_{av,i}$ and the unimodal feature $X_{vs,i}$ is formulated as:
@@ III-F Uncertainty-Driven Distillation
In practice, we design a network branch (a simple three-layer convolutional neural network, which is shown in Figure 3) to predict the variance ${\sigma}_{i}^{2}$ and reformulate the likelihood maximization problem as the minimization of a loss function. Specifically, we employ an uncertainty-weighted MSE loss:
@@
IV Experiments
IV-A Datasets and Evaluation Metrics
IV-A1 Datasets
@@
We conduct extensive experiments on two audio-visual benchmarks: XD-Violence [25] and CCTV-Fightssub [49], both of which contain synchronized audio and visual modalities. Unlike traditional unimodal datasets, these benchmarks enable a more comprehensive evaluation of our framework’s robustness under multimodal settings. XD-Violence. As the largest publicly available audio-visual WSVAD dataset, XD-Violence [25] significantly surpasses existing datasets in scale and diversity. It comprises 3,954 training videos and 800 test videos, with the test set containing 500 violent and 300 non-violent videos. The dataset covers six distinct categories of violent events, including abuse, car accident, explosion, fighting, riot, and shooting, which occur at various temporal locations within videos.
@@
CCTV-Fightssub. Derived from CCTV-Fights [77], CCTV-Fightssub [49] is a carefully curated subset designed to address audio-visual anomaly detection. The subset retains 644 high-quality videos depicting real-world fight scenarios, each with meaningful audio content, making it a valuable resource for evaluating audio-visual anomaly detection methods in real-world surveillance contexts.
IV-A2 Evaluation Metrics
@@
For performance evaluation, we adopt distinct metrics tailored to different granularities of WSVAD tasks. For coarse-grained WSVAD, we employ frame-level Average Precision (AP), which provides a comprehensive measure of detection accuracy across varying confidence thresholds. For fine-grained anomaly detection, we utilize mean Average Precision (mAP) [22] computed across multiple intersection over union (IoU) thresholds and the average mAP (AVG) across different thresholds. Specifically, we evaluate mAP at IoU thresholds ranging from 0.1 to 0.5 with an interval of 0.1, followed by reporting AVG across these thresholds.
IV-B Implementation Details
@@
We conduct experiments on an NVIDIA RTX 4090 GPU, where the visual enhancement network is a single-layer 1D convolutional network, which includes a convolutional layer with a kernel size of 3, padding size of 1, ReLU activation function, and a skip connection. Such an operation effectively facilitates the aggregation of local contextual information. For input processing, we employ a frame selection strategy tailored to different datasets, sampling one frame per 16 frames for XD-Violence and one frame per 4 frames for CCTV-Fightssub, using a uniform sampling strategy with a maximum frame count of 256; During optimization, we set the batch size, learning rate, and total epoch to 96, $1e^{-5}$, and 10, respectively.
TABLE I: Coarse-grained comparisons on XD-Violence.
@@
TABLE II: Coarse-grained comparisons on CCTV-Fightssub.
@@
IV-C Comparison with State-of-the-Art Methods
TABLE III: Fine-grained comparisons on XD-Violence.
Method Reference Modality mAP@IoU(%)
0.1 0.2 0.3 0.4 0.5 AVG
Random - RGB(VIT) 1.82 0.92 0.48 0.23 0.09 0.71
DeepMIL [12] CVPR 2018 RGB(ViT) 22.72 15.57 9.98 6.20 3.78 11.65
AVVD [22] TMM 2022 RGB(ViT) 30.51 25.75 20.18 14.83 9.79 20.21
VadCLIP [21] AAAI 2024 RGB(ViT) 37.03 30.84 23.38 17.90 14.31 24.70
AVadCLIP this work RGB(ViT) 39.63 32.77 26.84 21.58 16.39 27.44
AVadCLIP this work RGB(ViT)+Audio 41.89 34.61 27.08 22.16 17.30 28.61
TABLE IV: Fine-grained comparisons on CCTV-Fightssub.
Method Reference Modality mAP@IoU(%)
0.1 0.2 0.3 0.4 0.5 AVG
VadCLIP [21] AAAI 2024 RGB(ViT) 19.34 14.32 9.25 6.64 3.73 10.66
AVadCLIP this work RGB(ViT) 21.10 14.57 9.01 5.74 4.69 11.02
AVadCLIP this work RGB(ViT)+Audio 22.25 15.91 10.40 7.00 5.28 12.17
TABLE V: Cross-dataset WSVAD results on XD-Violence and CCTV-Fightssub.
[…]
XD-Violence 86.04 69.24
CCTV-Fightssub 76.60 73.38
IV-C1 Performance comparison on XD-Violence
Our experiments evaluate both coarse-grained and fine-grained anomaly detection performance on XD-Violence, comparing our AVadCLIP against state-of-the-art approaches, as shown in Tables IIII.
For coarse-grained anomaly detection, using only RGB inputs, AVadCLIP (* denotes RGB-only input) achieves an AP score of 85.53%, surpassing all existing vision-only methods. Notably, it outperforms VadCLIP, the previous best-performing RGB-only approach, by 1.0%, demonstrating superior visual anomaly detection. When incorporating audio, AVadCLIP further improves performance, significantly outperforming all multimodal baselines, achieving a remarkable 4.9% gain over the latest method AVCL [49].
For fine-grained anomaly detection, AVadCLIP consistently outperforms all competitors across different IoU thresholds, as detailed in Table III. With RGB-only input, AVadCLIP surpasses VadCLIP at all IoU thresholds, achieving an AVG improvement of 2.7%. Similarly, the full-modality model AVadCLIP leads across all metrics, boosting the AVG by 3.9%. These results highlight the effectiveness of multimodal learning in precisely localizing anomaly boundaries and improving category predictions.
IV-C2 Performance comparison on CCTV-Fightssub
The coarse-grained anomaly detection results on CCTV-Fightssub are presented in Table II. For RGB-only methods, AVadCLIP achieves 73.36% AP, surpassing the state-of-the-art VadCLIP and demonstrating the effectiveness of our approach in unimodal scenarios. For audio-visual scenarios, AVadCLIP further improves performance, outperforming all existing methods. These results indicate that incorporating audio information can further enhance anomaly detection performance, validating the effectiveness of cross-modal complementary information mining.
We present the fine-grained anomaly detection results on CCTV-Fightssub in Table IV, and it can be observed that AVadCLIP consistently outperforms all competitors at different IoU thresholds. Using RGB-only input, AVadCLIP surpasses VadCLIP at all IoU thresholds, achieving a 0.4% improvement in AVG. Similarly, AVadCLIP with audio leads in all metrics, increasing AVG by 1.5%. These results further highlight the effectiveness of multimodal learning in accurately locating anomalous boundaries and improving category prediction.
IV-C3 Cross-dataset Results
Table V presents the cross-dataset evaluation results of AVadCLIP on XD-Violence and CCTV-Fightssub, aiming to assess its generalization capability across different domains. Despite being trained on one dataset and tested on another, AVadCLIP consistently achieves competitive performance, demonstrating strong robustness and transferability. For example, AVadCLIP trained on XD-Violence still achieves an AP of 69.24% when directly tested on the surveillance-oriented CCTV-Fightssub, with less than a 4% drop compared to the model trained specifically on that dataset. These results highlight the model’s ability to generalize well to unseen data distributions and diverse anomaly scenarios.
Overall, AVadCLIP achieves state-of-the-art performance in both unimodal and multimodal settings across coarse-grained and fine-grained anomaly detection tasks. The comprehensive results validate its effectiveness in leveraging audio-visual collaboration and demonstrate the feasibility of uncertainty-driven distillation strategy.
IV-D Ablation Studies
TABLE VI: Effectiveness of designed modules on XD-Violence.